OpenStudy (anonymous):

Can someone help me make a Truth Table for this problem please.. ~p ^ (~r ^ q)

OpenStudy (amistre64):

i ored it ...

OpenStudy (amistre64):

set up r and q r -r q -r^q t f t f f t t t f f f t f f may be better

OpenStudy (amistre64):

p -p -p -r^q -p^(-r^q) t f f f f f t f t f f f f f f f t f f t t t t f f t f f i cant recall if "f.f = t" tho

OpenStudy (anonymous):

i have to show the whole process.. for the truth table.. my professor gave me this to start with p q r ~p ~r ~r ^ q ~p ^ (~r ^ q) T T T T T F T F T T F F F T T F T F F F T F F F now i just have to finish the rest.

OpenStudy (anonymous):

You understand the general method of constructing a truth table? \[\begin{array}{|ccc|c|c|c|c|c} p&q&r&\lnot p & \lnot r & \lnot r\wedge q & \lnot p \wedge(\lnot r \wedge q) \\ \hline T & T& T \\T & T & F \\ T&F & T \\ T & F & F \\ F & T& T \\ F & T & F \\ F & F & T \\ F & F & F\end{array}\]

OpenStudy (amistre64):

negate P negate r

OpenStudy (anonymous):

yes the ~ is for not... so it would be Not P

OpenStudy (anonymous):

Right, so now you just fill up the \(\lnot p\) column with the OPPOSITE of whatever p is for that row.

OpenStudy (amistre64):

id do -r^q first, but i think they are associative to begin with

OpenStudy (anonymous):

For example I'll do the first 5 in that column: \[\begin{array}{|ccc|c|c|c|c|c} p&q&r&\lnot p & \lnot r & \lnot r\wedge q & \lnot p \wedge(\lnot r \wedge q) \\ \hline T & T& T & F\\T & T & F & F\\ T&F & T & F \\ T & F & F & F \\ F & T& T & T\\ F & T & F \\ F & F & T \\ F & F & F\end{array}\]

OpenStudy (amistre64):

i get discrete math this term :) gonna be doing these in chapter 1

OpenStudy (anonymous):

Yep. I did a few of them last semester ;)

OpenStudy (anonymous):

So its the opposite for ~r too right?

OpenStudy (anonymous):

\(\lnot r \) is the opposite of r yes.

OpenStudy (anonymous):

Ok i got ~r done how

OpenStudy (anonymous):

how do i do the next row?

OpenStudy (anonymous):

Oh drat

OpenStudy (anonymous):

It's AND, not or.

OpenStudy (anonymous):

so the if either one have a T then its T?

OpenStudy (anonymous):

\(\lnot r\) must be T AND q must be T for \(\lnot r \wedge q\) to be T

OpenStudy (anonymous):

No, I got the operator wrong. \(\wedge\) is AND. I said OR originally. They must BOTH be True for this column to be True.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

practicing your tables Satellite? ;)

OpenStudy (anonymous):

so it should be F F F F T T F F

OpenStudy (anonymous):

\[\begin{array}{|ccc|c|c|c|c|c} p&q&r&\lnot p & \lnot r & \lnot r\wedge q & \lnot p \wedge(\lnot r \wedge q) \\ \hline T & T& T & F& F & F \\T & T & F & F& T & T \\ T&F & T & F & F & F \\ T & F & F & F &T & F \\ F & T& T & T & F & F \\ F & T & F &T & T & T \\ F & F & T &T & F & F \\ F & F & F&T & T & F \end{array}\]

OpenStudy (anonymous):

Is that what you got?

OpenStudy (anonymous):

yup thats what i got

OpenStudy (amistre64):

trying to leap over 3 columns gets my eyes lost :)

OpenStudy (anonymous):

lol mine too. so how do i do the final column?

OpenStudy (anonymous):

Ok so now the last column will be True if and only if the \(\lnot p\) is True and the \(\lnot r \wedge q\) column is True.

OpenStudy (anonymous):

Because it's \(\lnot p\) AND (\(\lnot r\) AND q)

OpenStudy (anonymous):

so it should be f f f f f t f f

OpenStudy (anonymous):

I honestly don't know why they use \(\wedge\) and \(\vee\) since it's not any harder to write AND or OR, and it's harder to mix them up.

OpenStudy (anonymous):

yea i agree with you there polnak

OpenStudy (amistre64):

dont be afraid to re order to columns to make life easier: \begin{array}{|ccc|c|c|c|c|c} p&r&q&\lnot r & \lnot r\wedge q & \lnot p & \lnot p \wedge(\lnot r \wedge q) \\ \end{array} perhaps

OpenStudy (anonymous):

mine is different

OpenStudy (anonymous):

That's because I screwed it up

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\begin{array}{|ccc|c|c|c|c|c} p&q&r&\lnot p & \lnot r & \lnot r\wedge q & \lnot p \wedge(\lnot r \wedge q) \\ \hline T & T& T & F& F & F & F \\T & T & F & F& T & T & F \\ T&F & T & F & F & F & F \\ T & F & F & F &T & F & F \\ F & T& T & T & F & F & F \\ F & T & F &T & T & T& T \\ F & F & T &T & F & F & F \\ F & F & F&T & T & F & F \end{array}\]

OpenStudy (anonymous):

Thats what i have :) :)

OpenStudy (amistre64):

p r q ---------- -F ^ -F ^ T should be the only "True" you get, regardless of the rest of it :)

OpenStudy (anonymous):

Thank you Polnak:) for helping me with this .. i'm sure i will be having MORE of these types of problems lol

OpenStudy (anonymous):

just keep your \(\wedge\) = AND, \(\vee\) = OR straight, and you'll be fine.

OpenStudy (anonymous):

ok thank you